# ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT

### v1

@@ 6.1. Failure cases analysis
To further understand the reasons that contribute to the failure of ZS4C in synthesizing compilable code even after applying Conversational Error Fixing, we conduct a qualitative analysis of the failed cases. We collect all 33 failed cases and manually examine the conversation history between the compiler and LLM. We derived the following five reasons.
@@ 6.3. Threats to validity
Threats to external validity. Consistent with prior studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), we utilize the StatType-SO benchmark as the evaluation dataset. This dataset, comprising manually verified code snippets from six prominent Java libraries across diverse domains, may not be representative of libraries in all programming languages. Nonetheless, ZS4C is language-agnostic and it does not rely on Java-specific constructs. We have chosen ChatGPT (GPT-3.5-Turbo) as the base LLM for our study. Consequently, our findings might differ if alternative LLMs, such as GPT-4, are employed.
Threats to internal validity. The primary threat to internal validity comes from the manually crafted prompt templates tailored for the task in ZS4C. We cannot guarantee the optimality of our prompts, nor can we consider all the potential prompts. To mitigate this threat, we elaborate on all of our design concerns and compare ZS4C with multiple baselines when basic prompts are utilized. Given the context window token constraints of ChatGPT, we consider adding up to one demonstration example to ZS4C and form its one-shot version, as incorporating more examples might breach the token limit. Furthermore, only the most recent chat history is retained in ConversationFixing once the context window limit is approached.

### v3

@@ 6. Discussion
6.1. Error fixing by ZS4C
To understand the errors that ZS4C is capable of fixing, we perform error analysis before and after running ZS4CGPT-3.5 on StatType-SO. To do so, we manually examine all the code snippets and their error logs returned by the compiler. We derived the following three types of errors for the code snippets of StatType-SO (Java code). Note that one code snippet could have multiple errors.
  • Symbol Not Found: It occurs during compilation when the compiler encounters an identifier (such as a variable or method name) that it cannot recognize or resolve.
  • Wrong Annotation: This error is triggered when an annotation is applied wrongly.
  • Method Override/Implementation Error: This error happens when a method signature does not match the one in the superclass or interface, violating the rules of method overriding or implementation.
We now show some examples of each type of error and how they were fixed by our approach.
In Figure 5, the first code snippet (5(a)) demonstrates “Symbol Not Found” errors. Specifically, the ConfigParameters class extends ParameterHolder and uses ResourceSettings and Environment classes without defining them. This causes multiple “Symbol Not Found” errors because these classes and their respective methods and fields are unrecognized. In the second code snippet (5(b)), these errors are resolved by defining the missing classes ParameterHolder, ResourceSettings, and Environment within the same file. This ensures that all referenced classes are recognized and accessible during compilation, thus eliminating errors.
public class xstream_class_34 {
@XStreamAlias("config")
public class ConfigParameters extends ParameterHolder {
/* Error: ParameterHolder
undefined */
@XStreamImplicit(itemFieldName = "resource")
private List<ResourceSettings> resources;
/* Error: ResourceSettings
undefined */
@XStreamImplicit(itemFieldName = "env")
private List<Environment> environments;
/* Error: Environment
undefined */
// Other code
(a) Code with errors (before running with ZS4C).
public class xstream_class_34 {
@XStreamAlias("config")
public static class ConfigParameters extends ParameterHolder {
/* Existing code (Figure a) */
/* Missing classes are added below */
public static class ParameterHolder { /* Fix: Added missing ParameterHolder class
definition */
public static class ResourceSettings { /* Fix: Added missing ResourceSettings class
definition */
private String name;
public String getName() { return name; }
public void setName(String name) { this.name = name; }
public static class Environment { /* Fix: Added missing Environment class
definition */
private String name;
public String getName() { return name; }
public void setName(String name) { this.name is name; }
(b) Code with fixes (after running with ZS4C).
Figure 5. An example of “Symbol Not Found” error fixes using ZS4C for StatType-SO.
In Figure 6, the first code snippet (6(a)) illustrates the “Wrong Annotation” errors. The annotation @Transactional is placed outside of any class, which causes the compiler to expect a class, interface, or enum declaration immediately following these annotations. This issue is addressed and fixed in the second code snippet (6(b)). The corrected placement of the annotations is directly before the GenericDaoImpl class inside the hibernate_class, ensuring proper association and adherence to Java syntax requirements.
package hibernate;
@Transactional
@SuppressWarnings("unchecked")
public class GenericDaoImpl<T, ID extends Serializable> implements GenericDao<T, ID> {
// Other code
(a) Code with errors (before running with ZS4C).
package hibernate;
public class hibernate_class { /* Fix: Correct annotation placement */
@Transactional
@SuppressWarnings("unchecked")
public class GenericDaoImpl<T, ID extends Serializable> implements GenericDao<T, ID> {
// Other code
(b) Code with fixes (after running with ZS4C).
Figure 6. An example of “Wrong Annotation” error fix using ZS4C for StatType-SO.
In Figure 7, the first code snippet (7(a)) shows “Method Override/Implementation Error”. The class DetailDollarsConverter extends the superclass ReflectionConverter but fails to correctly override the marshal and unmarshal methods because the method signatures do not match those in the superclass, causing the @Override annotations to throw errors. These issues are resolved in the second code snippet (7(b)) by ZS4C. The corrected class DetailDollarsConverter now implements the Converter interface, ensuring the correct implementation of the marshal and unmarshal methods by modifying the code structure and method definitions.
Table 9 presents the analysis performed on the StatType-SO dataset. The results indicate that ZS4C is highly effective in reducing various error types. This reduction is especially notable in the “Symbol Not Found” and “Method Override/Implement Error” categories, where every dataset showed a substantial decrease in errors. The results show the potential of ZS4C in other applications, such as fixing errors of code snippets that are generated by LLM and auto code completion in IDE (JetBrains Community, 2018; Pan et al., 2024).
public class xstream_class_24 {
public class DetailDollarsConverter extends ReflectionConverter {
@Override /* Error: Method does not override
marshal method in
ReflectionConverter */
public void marshal(Object obj, HierarchicalStreamWriter writer, MarshallingContext context) {
super.marshal(obj, writer, context);
writer.startNode("node4");
writer.setValue(Double.toString(20));
writer.endNode();
@Override /* Error: Method does not override
unmarshal method in
ReflectionConverter */
public Object unmarshal(HierarchicalStreamReader reader, UnmarshallingContext context) {
return super.unmarshal(reader, context);
// Other code
(a) Code with errors (before running with ZS4C).
public class xstream_class_24 {
public class DetailDollarsConverter implements Converter { /* Correct interface implementation */
private Mapper mapper;
private ReflectionProvider reflectionProvider;
public DetailDollarsConverter(Mapper mapper, ReflectionProvider reflectionProvider) {
this.mapper = mapper;
this.reflectionProvider = reflectionProvider;
@Override /* Fixed by implementing Converter
interface and defining marshal method
as outlined below */
public void marshal(Object obj, HierarchicalStreamWriter writer, MarshallingContext context) {
ReflectionConverter reflectionConverter = new ReflectionConverter(mapper, reflectionProvider);
reflectionConverter.marshal(obj, writer, context); /* Uses ReflectionConverter instance to
marshal the object */
writer.startNode("node4");
writer.setValue(Double.toString(20));
writer.endNode();
/* Following the above similar
logic, the ReflectionConverter instance
is utilized to unmarshal the object. */
// Other code
(b) Code with fixes (after running with our ZS4C).
Figure 7. Examples of “Method Override/Implementation” error fix using ZS4C for StatType-SO.
Table 9. Error analysis before and after running ZS4CGPT-3.5 on StatType-SO.
Before After
Symbol
Not Found
Wrong
Annotation
Method
Override/
Implementation
Error
Symbol
Not Found
Wrong
Annotation
Method
Override/
Implementation
Error
Android 50 0 21 2 0 0
JDK 21 0 3 0 0 0
GWT 50 0 9 4 0 0
Hibernate 50 1 2 17 0 1
Joda-Time 50 0 0 0 0 0
XStream 44 0 12 7 0 1
@@ 6.2. Failure cases analysis
To further understand the reasons that contribute to the failure of ZS4C with GPT-3.5 for StatType-SO dataset in synthesizing compilable code even after applying Conversational Error Fixing, we conduct a qualitative analysis of the failed cases. We collect all 33 failed cases and manually examine the conversation history between the compiler and LLM. We derived the following five reasons.
@@ 6.3. Limitations of ZS4C and potential solutions
6.4. Cost analysis of API calls
Table 10 summarizes the costs of ZS4C using the OpenAI API with GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4o) models along with StatType-SO and Python-SO datasets. The table compares the costs across one-shot, zero-shot, and three-shot prompt configurations for both models. As we can observe, it is economical to use both GPT-3.5 and GPT-4 to fix the code, which only costs 0.003 and 0.025 for one fix in the zero-shot setting, respectively. Using GPT-4 is more expensive than GPT-3.5, although it provides better performance. Meanwhile, the results of RQ4 indicate that zero-shot achieves comparable performance as one-shot and three-shot prompt settings. This finding is practically significant, as zero-shot prompt saves the cost of retrieving examples and reducing the prompt size, while still achieving comparable performance.
Table 10. Cost analysis of OpenAI API for various prompt settings based on input token usage.
StatType-SO Python-SO
Total Cost (USD) Cost/case (USD) Total Cost (USD) Cost/case (USD)
One-shot (GPT-3.5) \$1.143 \$0.004 \$1.635 \$0.003
Zero-shot (GPT-3.5) \$0.706 \$0.003 \$0.954, \$0.002
Three-shot (GPT-4) \$19.556 \$0.073 \$28.609 \$0.053
One-shot (GPT-4) \$10.92 \$0.041 \$15.826 \$0.029
Zero-shot (GPT-4) \$6.596 \$0.025 \$9.358 \$0.017
Total \$38.921 \$0.146 \$56.382 \$0.104
@@ 6.5. Threats to validity
Threats to external validity. Consistent with prior studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), we utilize the StatType-SO benchmark and a newly constructed Python dataset as the evaluation datasets. Those two datasets comprising manually verified code snippets across diverse domains, may not be representative of libraries in all programming languages. Nonetheless, ZS4C is language-agnostic and can be applied to any programming language. Another external threat is that the code snippets in our studied dataset are often small (with an average of 14-82 LOC across libraries), so our findings might not be generalized to larger code snippets. In this study, we have chosen GPT-3.5 and GPT-4 as our base LLMs for our study. Consequently, our findings might differ if alternative LLMs, such as Gemini, are employed. We encourage future research on more datasets with larger code snippets and more base LLMs.
Threats to internal validity. The primary threat to internal validity comes from the manually crafted prompt templates tailored for the task in ZS4C. We cannot guarantee the optimality of our prompts, nor can we consider all the potential prompts. To mitigate this threat, we elaborate on all of our design concerns and compare ZS4C with multiple baselines when basic prompts are utilized. Given the context window token constraints of ChatGPT, we consider adding up to one demonstration example to ZS4C and form its one-shot version, as incorporating more examples might breach the token limit. Furthermore, only the most recent chat history is retained in ConversationFixing once the context window limit is approached for GPT-3.5.
As introduced in Section 3.2, the same APIs could appear in different packages (i.e., API collision) and make it challenging to decide which packages to use during Conversational Error Fixing. This could be an internal threat to the validity. To mitigate the threat, we selected the one that satisfies pre-defined constraints (i.e., invoked functions and attributes) following the previous study (Subramanian et al., 2014). If there exist multiple libraries available after resolving constraints, we select the one that was released most recently. Through such process, we achieved a low API collision rate 5.5% (82 out of 1,494 cases exhibit API collisions) on StatType-SO dataset.